OpenStudy (anonymous):

what is the length on MN? ML=3x MN= x+8 LN=?

OpenStudy (anonymous):

L is the midpoint right?

OpenStudy (anonymous):

its a triangle

OpenStudy (anonymous):

is there a picture of it?

OpenStudy (anonymous):

no but its an equilateral triangle

OpenStudy (anonymous):

ML=MN 3x=x+8 2x=8 x=4 MN: x+8 4+8=12 ML: 3(4)=12

OpenStudy (anonymous):

Since the triangle is equilateral, the length of LN will also be 12

OpenStudy (anonymous):

oh duh haha thank you

OpenStudy (anonymous):

np

OpenStudy (anonymous):

thx